a truck enters the highway driving 60mph. a car enter the highway at the same place 8 minutes later and drives 67mph in the same direction. from the time the car enters the highway , how long will it take the car to pass the truck?

Answer by rfer(16322)

8 minutes = 8/60=0.1333333 hrs
60*0.133333=8 mi lead
overtake at 7mph
8/7=1.14 hrs or 1hr 8.5 min